# Graph Neural Network-based EEG Classification: A Survey

### v1

@@ II Overview of Graph Neural Networks
Multiple types of GNNs have been well introduced in wu_comprehensive_2021 ; zhou_graph_2020 . In this survey, we briefly introduce the two main branches of GNNs, namely, spatial and spectral GNNs (Fig. 2). Other types of GNNs, such as attention GNNs velickovic_gat_2018 , recurrent GNNs seo_structured_2018 , and graph transformers shi_masked_2021 , can be viewed as special cases of spatial GNNs, and thus we will not provide detailed discussion in this survey. Both spatial and spectral GNNs aim to extend the convolution mechanism to graph data. For a detailed review of their similarities and differences, see chen_bridging_2021 .
@@
Figure 2: Illustration of core mechanisms of spatial and spectral graph neural networks (GNNs). A) An undirected featured graph is given as an example input graph with node features shown as node labels and colours. B) Spatial GNNs operate in the graph domain directly using message passing to update node embeddings. 1) Messages, i.e. transformed node features or embeddings, are sent along edges. For simplicity, we show only one direction of the flow of messages. 2) The collected messages at each node are aggregated using a permutation-invariant function and are fused with the original node embedding to form an updated node embedding. Thus, one spatial GNN layer results in node embeddings containing information about the 1-hop neighbourhood of a given node. Thus, $L$ layers are required for node embeddings to access the information from the $L$-hop neighbourhood. C) In contrast, spectral GNNs operate in the graph spectral domain. 1) Node features are treated as signals on top of a graph and are deconstructed into graph frequencies given by the eigendecomposition of the graph Laplacian. Graph frequencies can be interpreted as variations of the signals. 2) The contribution of each graph frequency is weighted by the set of learnable kernels $G$ that effectively function as graph filters. 3) Node embeddings are then obtained by aggregating the filtered graph frequencies and transforming them back to the spatial graph domain. Thus, full spectral GNNs can access information from $N$-hop neighbourhoods where $N$ is the number of nodes of a given graph. However, in practice, approximations such as Chebyshev graph convolution restrict this to the chosen hop size.
@@ II.1 Spatial GNNs
where $h_{i}$ is the node embedding vector, or when $l=1$, this is the input node feature vector. $\sigma$ is the activation function, $\sum$ is the aggregation function, $\mathcal{N}(v_{i})$ is the neighbourhood of node $v_{i}$, $W\in\mathbb{R}^{d_{1}\times d_{2}}$ is a learnable parameter matrix projecting node embeddings from input dimension $d_{1}$ to hidden dimension $d_{2}$ and $e_{ji}$ is the edge weight ($e_{ji}=1$ for unweighted graphs).
@@ II.2 Spectral GNNs
The graph spectrum is defined as the eigendecomposition of the graph Laplacian matrix. The GFT is then defined as $\mathbf{\hat{H}}=\mathbf{U}^{T}\mathbf{H}$, its inverse as $\mathbf{H}=\mathbf{U}\mathbf{\hat{H}}$, where $\mathbf{U}$ is the orthonormal matrix of eigenvectors of the graph Laplacian $\mathbf{L}$ and $H\in\mathbf{R}^{N\times D}$ is the matrix of node feature vectors with $N$ and $D$ being the number of nodes and dimensionality of node features, respectively. The graph Laplacian is defined as $\mathbf{L}=D-A$, but often the normalized version is preferred: $\mathbf{\hat{L}}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ ($\mathbf{A}$ and $\mathbf{D}$ are the adjacency and degree matrices, respectively).

### v2

@@ II Overview of Graph Neural Networks
Compared to other deep learning models, GNNs offer several advantages. First, GNNs were specifically designed for graph-structured inputs. This means that GNNs can adapt to irregularly structured inputs, i.e. graphs with varying numbers of nodes, compared to traditional deep learning, such as CNN, that require fixed-size inputs. Next, GNNs can simultaneously learn information from node features and the graph structure by accepting two inputs: node feature matrix and graph structure. Such simultaneous integration is not possible with traditional deep learning methods.
Multiple types of GNNs have been well introduced in wu_comprehensive_2021 ; zhou_graph_2020 . In this survey, we briefly introduce the two main branches of GNNs, namely, spatial and spectral GNNs (Fig. 2). Other types of GNNs, such as attention GNNs velickovic_gat_2018 , recurrent GNNs seo_structured_2018 , and graph transformers shi_masked_2021 , can be viewed as special cases of spatial GNNs, and thus we will not provide detailed discussion in this survey. Both spatial and spectral GNNs aim to extend the convolution mechanism to graph data. For a detailed review of their similarities and differences, see chen_bridging_2021 . Moreover, for a comparison of GNNs in terms of computational complexity, see wu_comprehensive_2021 .
@@ II.1 Spatial GNNs
Figure 2: Illustration of core mechanisms of spatial and spectral GNNs. A) An undirected featured graph is given as an example input graph with node features shown as node labels and colours. B) Spatial GNNs operate in the graph domain directly using message passing to update node embeddings. 1) Messages, i.e. transformed node features or embeddings, are sent along edges. For simplicity, we show only one direction of the flow of messages. 2) The collected messages at each node are aggregated using a permutation-invariant function and are fused with the original node embedding to form an updated node embedding. Thus, one spatial GNN layer results in node embeddings containing information about the 1-hop neighbourhood of a given node. Thus, $L$ layers are required for node embeddings to access the information from the $L$-hop neighbourhood. C) In contrast, spectral GNNs operate in the graph spectral domain. 1) Node features are treated as signals on top of a graph and are deconstructed into graph frequencies given by the eigendecomposition of the graph Laplacian. Graph frequencies can be interpreted as variations of the signals. 2) The contribution of each graph frequency is weighted by the set of learnable kernels $G$ that effectively function as graph filters. 3) Node embeddings are then obtained by aggregating the filtered graph frequencies and transforming them back to the spatial graph domain. Thus, full spectral GNNs can access information from $N$-hop neighbourhoods where $N$ is the number of nodes of a given graph. However, in practice, approximations such as Chebyshev graph convolution restrict this to the chosen hop size.
@@
where $h_{i}$ is the node embedding vector, or when $l=1$, this is the input node feature vector, $\sigma$ is the activation function, $\sum$ is the aggregation function, $\mathcal{N}(v_{i})$ is the neighbourhood of node $v_{i}$, $W\in\mathbb{R}^{d_{1}\times d_{2}}$ is a learnable parameter matrix projecting node embeddings from input dimension $d_{1}$ to hidden dimension $d_{2}$ and $e_{ji}$ is the edge weight ($e_{ji}=1$ for unweighted graphs).
@@ II.2 Spectral GNNs
The graph spectrum is defined as the eigendecomposition of the graph Laplacian matrix. The GFT is then defined as $\mathbf{\hat{H}}=\mathbf{U}^{T}\mathbf{H}$, its inverse as $\mathbf{H}=\mathbf{U}\mathbf{\hat{H}}$, where $\mathbf{U}$ is the orthonormal matrix of eigenvectors of the graph Laplacian $\mathbf{L}$ and $H\in\mathbf{R}^{N\times D}$ is the matrix of node feature vectors with $N$ and $D$ being the number of nodes and dimensionality of node features, respectively. The graph Laplacian is defined as $\mathbf{L}=\mathbf{D}-\mathbf{A}$, but often the normalised version is preferred: $\mathbf{\hat{L}}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ ($\mathbf{A}$ and $\mathbf{D}$ are the adjacency and degree matrices, respectively).